tutors please help me find the vertex and focus of the equatin y^2 = -8



Answer by DSMLMD(16)

I can't find the vertex and focus point of the equation y^2 = -8. It's a quadratic equation, but if the equation are y^2 = -8x, there's able to find the vertex and focus point.


y^2 = -8x
y^2 = 4(-2)x


because y^2 = -8x is a parabola equation y^2 = 4px. So we get:


vertex:
(0,0)


focus point:
(p,0)
(-2,0)